OpenStudy (anonymous):

a cube is circumscribed by a sphere with a 1-diameter. how much more volume is the sphere

OpenStudy (anonymous):

Vol of Sphere = 4 pi/3 Vol of cube = 2 root 2 4pi/3 - 2 root 2 = 4.189 - 2.828 = -1.36 approx

OpenStudy (anonymous):

Sorry, 1.36